# Explaining time series models using frequency masking

### v1

@@ 4.1 Synthetic data results
Table 1 shows the localization and complexity scores across the different methods for both the model trained on the low noise and the noisy datasets. The results on the low noise model show that all XAI methods perform approximately equal on the localization score, while IG and LRP yield relevance maps with substantially lower complexity scores compared to FreqRISE. However, when we move to the noisy model, the localization scorre of FreqRISE is unchanged, whereas both IG and LRP have much lower performance. The complexity of the IG and LRP relevance maps is slightly higher on the noisy model, while the complexity of the FreqRISE is unchanged. An example of the computed relevance maps is shown in the appendix.
Table 1: Localization (L) and complexity (C) on the synthetic data.
@@ 4.2 AudioMNIST results
Following the procedure described in Section 3.4, we initially compute the faithfulness results. Figure 2 shows the deletion plots for the digit classification task in the frequency and time-frequency domain. We have also included two additional baselines, namely randomly deleting features and deleting features based on their amplitude. In both domains, the mean true class probability quickly drops. After dropping only 15% of the features, FreqRISE has a mean true class probility of $0.206$ in the time-frequency domain and $0.194$ in the frequency domain. After this, the value continues to drop with features being removed.
Figure 3 shows the same results for the gender classification task. Here, the mean true class probability using FreqRISE is $0.434$ after dropping only 5% of the features in the frequency domain, while the same value in the time-frequency domain is $0.836$. The other methods show a different trend in time-frequency domain, but the results show that FreqRISE has a more difficult time masking out relevant features in the time-frequency domain.
Figure 2: AudioMNIST: Deletion plots for the digit task.
Figure 3: AudioMNIST: Deletion plots for the gender task.
Table 2: AudioMNIST: Faithfulness of explanations in the time (T), frequency (F) and time-frequency (TF) domain.
@@
IG .$183$$251$$194$$470$$428$$\mathbf{.389}$
LRP .$\mathbf{.168}$$204$$209$$401$$431$420.420.420.420
(Freq)RISE .$212$$\mathbf{.147}$$\mathbf{.125}$$\mathbf{.399}$$\mathbf{.414}$525.525.525.525
In Table 2, the faithfulness scores are shown for both models in all three domains. We notice that in the time domain, RISE performs either worse or equivalently compared to LRP. However, when we move to the frequency domain, the faithfulness scores are lowest for FreqRISE in both cases. Finally, when we look at the time-frequency domain, the faithfulness score increases for FreqRISE on the gender task, while it decreases on the digit task. These results indicate that the information is sparser for the digit task in the time-frequency domain.
Table 3 shows the complexity scores across all methods, domains and tasks. Again, it is clear that the complexity of the masking based models is higher compared to the remaining methods.
Table 3: AudioMNIST: Complexity scores for explanations in the time (T), frequency (F) and time-frequency (TF) domains.
@@
(a) AudioMNIST: Gender classification, relevance map in the time domain.
(b) AudioMNIST: Gender classification, relevance map in the frequency domain.
Figure 5: AudioMNIST: Digit classification, relevance map in the time-frequency domain.
4(a) shows an example of relevance maps computed in the time domain on the gender classification task. 4(b) shows the same sample in the frequency domain. The sample is from a female speaker and the model correctly predicts the class. We see that LRP has a very sparse but scattered signal in the time domain, while RISE appears very noisy. Moving to the frequency domain, the relevance is much more localized, with FreqRISE putting most emphasis on the fundamental frequency, while LRP also focuses on the harmonics. The fundamental frequency is known to be discriminative of gender (Baken & Orlikoff, 2000).
Figure 5 shows the relevance map computed using Freq-RISE in the time-frequency domain for the digit task. The spoken and predicted digit is 9. The relevance map shows that most relevance is put just at the beginning of the signal. However, around $t=0.5$, there are two distinct patterns emerging along the frequency axis. This shows the benefit of having both the time and frequency component when computing relevance maps for the digit task.
@@ 5 Discussion and conclusion
Time series data is inherently difficult to interpret, due to the complex patterns of the signals. Therefore, providing relevance maps in a domain where information is sparser and more easily interpretable is desirable. We therefore proposed FreqRISE, which computes relevance maps in the frequency and time-frequency domains using model-agnostic masking methods. Vielhaben et al. (2024) previously produced relevance maps in these domains using gradient based methods, which are model-dependent, limiting the usability in cases where model gradients are not available.
RISE and FreqRISE consistently yield high complexity scores. While LRP and IG are able to assign zero relevance, this is generally not the case for Freq(RISE). This is, however, not necessarily indicative of the perceptual quality of the relevance maps. The entropy could e.g. be reduced through postprocessing of relevance maps.
@@
We used a synthetic dataset with known salient features in the frequency domain to compute localization scores. In the low noise setting, FreqRISE gave slightly higher accuracy in identifying the correct frequency components. However, when testing in the high noise setting, FreqRISE outperformed the two baseline methods by a large margin (100% vs. 64%). These results indicate that FreqRISE is more robust in correctly identifying salient features in the frequency domain.
On AudioMNIST, we measured the faithfulness of all methods across three domains. RISE performed either slightly worse or similar to the baseline methods in the time domain. However, in the frequency domain FreqRISE gave the best faithfulness scores. In the time-frequency domain, FreqRISE gave the best performance on the digit task, but the worst performance on the gender task. This leads us to believe that FreqRISE is most suitable for providing explanations in domains where the information is assumed to be sparse. While the same information can theoretically be found in e.g. the time domain, the intricacy of the patterns would require more advanced masking schemes to identify the same patterns.
We believe that XAI methods that can provide frequency and time-frequency relevance maps are important components in ensuring safe and trustworthy time series models.
Impact Statement
This paper presents work whose goal is to advance the field of explainable AI. Explainable AI is a necessary tool in obtaining safe and trustworthy models. However, if the explanations are not correct, they can potentially create a false sense of confidence in the models and cause errors in critical domains.
Appendix A Design choices for FreqRISE
@@
For AudioMNIST, FreqRISE needs a large number of masks to converge ($N=20,000$). We found that when using fewer masks, the method still finds the relevant bits, but sorting out the irrelevant bits of the signal requires more masks. Wickstrøm et al. (2023) presented theoretical results on computing the number of masks and found that $N=3,000$ should yield results with a low error. It would be interesting to investigate convergence properties for our datasets in light of the theoretical results.
Additionally, Mercier et al. (2022) propose TimeREISE which uses a multiple mask sizes and sampling probabilities to compute the relevance maps for each time series. Finally, Cooper et al. (2022) apply RISE to images and propose a hierarchical systematic mapping to reduce the number of masks. Both of these avenues could be interesting to explore when masking in the frequency domain.
Appendix B Results on synthetic data

### v2

@@ 4 Results
We present a comprehensive evaluation of FreqRISE using the experimental setup described in the previous section.
@@ 4.1 Synthetic data results
Table 1 shows the localization and complexity scores across the different methods for both the model trained on the low noise and the noisy datasets. The results on the low noise model show that all XAI methods perform approximately equal on the localization score, while IG and LRP yield relevance maps with substantially lower complexity scores compared to the FreqRISE relevance maps. However, when we move to the noisy model, the localization score of FreqRISE is unchanged, whereas both IG and LRP have much lower performance. The unprocessed FreqRISE relevance maps generally have a high complexity compared to both IG and LRP. However, when post-processing the relevance maps at $p=0.997$, FreqRISE achieves complexity scores lower than or comparable to both baselines. At the same time, the localization performance of the post-processed relevance maps remains unchanged still outperforming the baselines.
Table 1: Localization (L) and complexity (C) on the synthetic data. FreqRISEp refers to the relevance maps post-processed with $p=0.997$.
@@ 4.2 AudioMNIST results
Subsequently, we compute results on the AudioMNIST tasks in the time, frequency and time-frequency domain.
Following the procedure described in Section 3.4, we initially compute the faithfulness results. Figure 3 (top) shows the deletion plots for the digit classification task in the frequency and time-frequency domain. We have also included two additional baselines, namely randomly deleting features (Rand.) and deleting features based on their amplitude (Amp.). In both domains, the mean true class probability quickly drops. After dropping only 10% of the features, FreqRISE has a mean true class probability of $0.139$ in the time-frequency domain and $0.248$ in the frequency domain. After this, the value continues to drop with features being removed.
Figure 3 (bottom) shows the same results for the gender classification task. Here, the mean true class probability using FreqRISE is $0.435$ after dropping only 5% of the features in the frequency domain, and the same value in the time-frequency domain is $0.445$. The other methods drop to lower values in the time-frequency domain, indicating that FreqRISE struggles more to identify relevant features in the time-frequency domain.
Figure 3: AudioMNIST: Deletion plots for both tasks in frequency and time-frequency domains. We delete features according to the importance and measure model outputs. FreqRISE outperforms the baselines in both domains on the digit task and the frequency domain on the gender task.
In Table 2 (top), the faithfulness scores are shown for both models in all three domains. We notice that in the time domain, RISE performs either worse or equivalently compared to LRP. However, when we move to the frequency domain, the faithfulness scores are lowest for FreqRISE in both cases. Finally, when we look at the time-frequency domain, the faithfulness score increases for FreqRISE on the gender task, while it decreases on the digit task. These results indicate that the information is sparser for the digit task in the time-frequency domain.
Table 2 (bottom) shows the complexity scores across all methods, domains and tasks. It is clear that the complexities of RISE and FreqRISE are higher compared to IG and LRP.
Table 2: AudioMNIST: Faithfulness (top) and Complexity (bottom) scores for explanations in the time (T), frequency (F) and time-frequency (TF) domains.
@@
IG .$\mathbf{.121}$$150$$127$$250$$232$$\mathbf{.212}$
LRP .$144$$128$$134$$\mathbf{.231}$$235$228.228.228.228
(Freq)RISE .$207$$\mathbf{.119}$$\mathbf{.076}$$266$$\mathbf{.226}$230.230.230.230
@@
Figure 4: AudioMNIST: Gender task. We show relevance maps computed by LRP and FreqRISE in the time and frequency domain. The salient information is more localized in the frequency domain.
Figure 5: AudioMNIST: Digit task. STDFT of the signal on the left and the FreqRISE relevance map in the time-frequency domain on the right. The relevance map shows the benefit of having both the time and frequency axis, when the describing the digit data.
Figure 4 (top) shows an example of relevance maps computed in the time domain on the gender classification task. Figure 4 (bottom) shows the same sample in the frequency domain. The sample is from a female speaker and the model correctly predicts the class. We see that LRP has a very sparse but scattered signal in the time domain, while RISE appears very noisy. Moving to the frequency domain, the relevance is much more localized, with FreqRISE putting most emphasis on the fundamental frequency, while LRP also focuses on the harmonics. The fundamental frequency is known to be discriminative of gender [baken2000clinical].
Figure 5 shows the relevance map computed using FreqRISE in the time-frequency domain for the digit task. The spoken and predicted digit is 9. The STDFT reveals that the signal starts around $t=0.3s$, where the relevance map puts most of the relevance. However, around $t=0.5s$, there are two distinct patterns emerging along the frequency axis. The fact, that the method is allowed to distribute the relevance differently across frequencies depending on the time, shows the benefit of having both the time and frequency components when computing relevance maps for the digit task.
4.2.1 Post-processing results
Here, we present faithfulness and complexity scores when post-processing the AudioMNIST relevance maps in the frequency and time-frequency domains.
Figure 6 shows complexity and faithfulness scores for different $p$-values for the post-processing of FreqRISE relevance maps as well as for LRP and IG. Ideally, both faithfulness and complexity scores should be low. Thus, an ideal explanation method would give results in the lower, left quadrant of the plot. From Figure 6 it is clear that with existing methods, faithfulness and complexity are typically trade-offs of each other.
For the digit task, FreqRISE consistently gives better faithfulness scores than the baselines for almost all values of $p$ in both domains. In the frequency domain, it is also clear that we are able to reduce the complexity scores to a level close to the LRP complexity, while maintaining a better faithfulness score (at $p=0.8$). In the time-frequency domain, we can similarly reduce the complexity, although not to a competitive level. However, here the gap in faithfulness between FreqRISE and competing methods is quite substantial, indicating that maybe a higher complexity is needed to properly describe the data.
For the gender task in the frequency domain, we still achieve better faithfulness scores for all thresholds. Additionally, the best complexity score of FreqRISE is comparable to both IG and LRP. In the time-frequency domain FreqRISE generally performs worse on both faithfulness and complexity. This is in line with the results presented.
Generally, we see that with simple thresholding of the FreqRISE relevance maps, we are able to reduce complexities of FreqRISE relevance maps while keeping competitive faithfulness scores.
Figure 6: Post-processing results at different thresholds for AudioMNIST relevance maps.
@@ 5 Discussion and conclusion
Time series data is inherently difficult to interpret, due to the complex patterns of the signals [rojat2021a]. Therefore, providing relevance maps in a domain where information is sparser and more easily interpretable is desirable. We therefore proposed FreqRISE, which computes relevance maps in the frequency and time-frequency domains using model-agnostic masking methods. vilrp previously produced relevance maps in these domains using gradient based methods, which are model-dependent, limiting the usability in cases where model gradients are not available.
@@
We used a synthetic dataset with known salient features in the frequency domain to compute localization scores. In the low noise setting, FreqRISE gave slightly higher accuracy in identifying the correct frequency components. However, when testing in the high noise setting, FreqRISE outperformed the two baseline methods by a large margin (100% vs. 64%). These results indicate that the gradient-based methods (IG and LRP) both are susceptible to noise in the data, resulting in a lower localization score in the noisy setting. Clearly, FreqRISE does not suffer from the same issues and instead has a stable, high performance independent of the noise level. Likely because it is not dependent on the gradients of the model.
On AudioMNIST, we measured the faithfulness of all methods across three domains. RISE performed either slightly worse or similar to the baseline methods in the time domain. However, in the frequency domain, FreqRISE gave the best faithfulness scores. In the time-frequency domain, FreqRISE gave the best performance on the digit task, but the worst performance on the gender task. The difference is likely to be found in the different data properties of the two tasks. For the gender task, the frequency domain is sufficient for describing the data, since the fundamental frequency is known to be discriminative of gender baken2000clinical. Thus, when moving to the time-frequency domain, unnecessary complexity is added to the problem, resulting in a lower performance by FreqRISE. This leads us to believe that FreqRISE gives the best performance when computed in a domain where the salient information is sparse. As of now, no existing methods are able to estimate the ”correct” domain in which to explain a model. It is therefore up to the domain experts to choose which domain is more meaningful to look at. Future work should focus on developing metrics that measure the most informative domains.
RISE and FreqRISE consistently yield high complexity scores. While LRP and IG are able to assign zero relevance, this is generally not the case for Freq(RISE). This is, however, not necessarily indicative of the perceptual quality of the relevance maps. We show that through simple post-processing, we can in many cases lower the complexity values to competitive levels while keeping a better faithfulness score (in the AudioMNIST case) or localization score (in the synthetic data case) than competing baselines. Future work should focus on a more rigorous heuristic for choosing suitable thresholds for post-processing of the FreqRISE relevance maps.
Like many other explainabilty methods, FreqRISE requires choosing a suitable set of hyper-parameters. This includes choosing an appropriate size of the lower dimensional grid on which we sample the binary masks, as well as the number of masks to use. As of now, we have no principled way of choosing these hyperparameters, although one solution could be to tune the faithfulness and complexity scores on a small validation set. Another, less systematic approach, simply requires the user to estimate the granularity of the problem and choose the grid size based on this. Finally, FreqRISE requires a rather large number of masks ($N=10,000$) in order to produce high-quality relevance maps for AudioMNIST. This clearly increases the computational complexity compared to competing methods. Future work should investigate how to monitor convergence for potential early stopping.
In conclusion, FreqRISE gives stable, good performance in localization and faithfulness on two different datasets with a total of 4 different tasks. Additionally, unlike existing methods, FreqRISE is completely model-independent requiring no access to gradients or weights of the models being explained. As such, FreqRISE is an important next step in providing robust explanations in the frequency and time-frequency domain for time series data.
[…]
Appendix A Appendix
A.1 Details on the math behind FreqRISE
We closely follow the approach presented in rise, but extend the notation to include FreqRISE, where we allow for masking in dual domains. First, we reiterate the notation introduced in Section 2.1.
We consider a time series input $\boldsymbol{X}\in\mathbb{R}^{V\times T}$ transformed into some domain, $S$, through the invertible transform $g:\boldsymbol{X}\rightarrow\boldsymbol{X}^{S}$. Generally, $\boldsymbol{X}^{S}=\left\{X^{S}\mid X^{S}:\Lambda\rightarrow\mathbb{R}^{3}\right\}$ is of size $V\times T_{S}\times F$, i.e. $\Lambda=\{1,\dots,V\}\times\{1,\dots,T_{S}\}\times\{1,\dots,F\}$. If $g$ is the Discrete Fourier transform $T_{S}=1$, while if $g$ is the identity (i.e. we stay in the time domain) $F=1$ and $T_{S}=T$. To ease notation, we will assume $V=1$, but the math can easily be extended to the case where $V>1$.
We then sample binary masks $\boldsymbol{M}:\Lambda\rightarrow\{0,1\}$ from distribution $\mathcal{D}$. The perturbed input $\hat{\boldsymbol{X}}\left(\boldsymbol{M}\right)$ is then obtained through elementwise multiplication with mask $\boldsymbol{M}$ in domain $S$ and subsequent inverse transformation through $g^{-1}$ back to the time domain:
$\hat{\boldsymbol{X}}\left(\boldsymbol{M}\right)=g^{-1}\left(\boldsymbol{X}^{S}\odot\boldsymbol{M}\right).$ (7)
We then obtain the outputs $\hat{y}_{c}\left(\boldsymbol{M}\right)$ as the class prediction for class, $c$, using the black box model $f$:
$\hat{y}_{c}\left(\boldsymbol{M}\right)=f\left(\hat{\boldsymbol{X}}\left(\boldsymbol{M}\right)\right)_{c}.$ (8)
Let us now define the importance, $\boldsymbol{R}_{c}\left(\lambda\right)$, of element $\lambda=(t,f)$ as the expected value over all possible masks $\boldsymbol{M}$, conditioned on $\lambda$ being observed, i.e. $\boldsymbol{M}(\lambda)=1$:
$\boldsymbol{R}_{c}\left(\lambda\right)=\mathbb{E}_{\boldsymbol{M}}\left[\hat{y}_{c}\left(\boldsymbol{M}\right)\mid\boldsymbol{M}(\lambda)=1\right].$ (9)
Since $\boldsymbol{M}$ is binary, the expectation can be computed as a sum over masks, $\boldsymbol{M}^{\prime}$:
$\boldsymbol{R}_{c}\left(\lambda\right)=\sum_{\boldsymbol{M}^{\prime}}\hat{y}_{c}\left(\boldsymbol{M}^{\prime}\right)P\left[\boldsymbol{M}=\boldsymbol{M}^{\prime}\mid\boldsymbol{M}\left(\lambda\right)=1\right].$ (10)
Using the definition of conditional distributions, this can be rewritten as:
$\boldsymbol{R}_{c}\left(\lambda\right)=\frac{1}{P\left[\boldsymbol{M}\left(\lambda\right)=1\right]}\sum_{\boldsymbol{M}^{\prime}}\hat{y}_{c}\left(\boldsymbol{M}^{\prime}\right)P\left[\boldsymbol{M}=\boldsymbol{M}^{\prime},\boldsymbol{M}\left(\lambda\right)=1\right].$ (11)
Now, again since $\boldsymbol{M}(\lambda)$ is binary,
$P\left[\boldsymbol{M}=\boldsymbol{M}^{\prime},\boldsymbol{M}\left(\lambda\right)=1\right]$ =$=\begin{cases}0,&\text{if $\boldsymbol{M}^{\prime}\left(\lambda\right)=0$.}\\ P\left[\boldsymbol{M}=\boldsymbol{M}^{\prime}\right],&\text{if $\boldsymbol{M}^{\prime}\left(\lambda\right)=1$.}\end{cases}$ (12)
$=\boldsymbol{M}^{\prime}\left(\lambda\right)\cdot P\left[\boldsymbol{M}=\boldsymbol{M}^{\prime}\right].$
Additionally, $P\left[\boldsymbol{M}\left(\lambda\right)=1\right]=\mathbb{E}\left[\boldsymbol{M}\left(\lambda\right)\right]$. Substituting this into (11):
$\boldsymbol{R}_{c}\left(\lambda\right)=\frac{1}{\mathbb{E}\left[\boldsymbol{M}\left(\lambda\right)\right]}\sum_{\boldsymbol{M}^{\prime}}\hat{y}_{c}\left(\boldsymbol{M}^{\prime}\right)\cdot\boldsymbol{M}^{\prime}\left(\lambda\right)\cdot P\left[\boldsymbol{M}=\boldsymbol{M}^{\prime}\right].$ (13)
We can now write this in matrix notation, such that the relevance can be computed for all elements, $\lambda$, simultaneously:
$\boldsymbol{R}_{c}=\frac{1}{\mathbb{E}\left[\boldsymbol{M}\right]}\sum_{\boldsymbol{M}^{\prime}}\hat{y}_{c}\left(\boldsymbol{M}^{\prime}\right)\cdot\boldsymbol{M}^{\prime}\cdot P\left[\boldsymbol{M}=\boldsymbol{M}^{\prime}\right].$ (14)
The relevance in (14) is then empirically estimated through Monte Carlo sampling. We sample $N$ masks, $\{\boldsymbol{M}_{1},\dots,\boldsymbol{M}_{N}\}$ according to $\mathcal{D}$ and normalize by the expectation of $\boldsymbol{M}$:
$\boldsymbol{R}_{c}\approx\frac{1}{N\cdot\mathbb{E}[\boldsymbol{M}]}\sum_{n=1}^{N}\hat{y}_{c}\left(\boldsymbol{M}_{n}\right)\cdot\boldsymbol{M}_{n},$ (15)
finally arriving at Equation (3) in the main paper.
A.2 Design choices for FreqRISE
@@
For AudioMNIST, FreqRISE needs a large number of masks to converge ($N=10,000$). We found that when using fewer masks, the method still finds the relevant bits, but sorting out the irrelevant bits of the signal requires more masks. relax presented theoretical results on computing the number of masks and found that $N=3,000$ should yield results with a low error. It would be interesting to investigate convergence properties for our datasets in light of the theoretical results.
Additionally, mercier2022a propose TimeREISE which uses a multiple mask sizes and sampling probabilities to compute the relevance maps for each time series. Finally, hipe apply RISE to images and propose a hierarchical systematic mapping to reduce the number of masks. Both of these avenues could be interesting to explore when masking in the frequency domain.
A.3 Additional relevance maps on AudioMNIST
Additional relevance maps in all domains on the digit and gender task for AudioMNIST.
Figure 7 shows the same example as Figure 5 in the paper, but also in time and frequency. 7(a) shows that LRP seems to highlight most of the STDFT as important, seemingly ranked by the amplitude of the actual signal. This is not very useful for explanations. The time domain plots show that the time relevance of LRP seems to be centered around the middle of the signal. While for RISE, the time relevance is still scattered. In the frequency domain, the information seems more localized for both LRP and FreqRISE.
Figure 8 shows an example on the digit task with the digit 0 being spoken. Here, we again see a distinct pattern in time and frequency for FreqRISE, while LRP seemingly focuses mostly on the lowest frequency component. In the time domain, the relevance is again scattered for both methods. In the frequency domain, the relevance is mostly localized on two frequency components for FreqRISE.
Figure 9 shows the same example as Figure 4 in the main paper, but also in the STDFT domain. In the STDFT domain, we see that FreqRISE struggles to identify relevant information, whereas LRP distributes relevance on the three lowest frequency components. This indicates that the frequency domain is sufficient for explaining the gender task.
Figure 10 shows an example on the gender task for a male speaker. In the STDFT domain, FreqRISE seemingly just highlights most of the signal, while LRP again puts emphasize on the three lowest frequency components. In the time domain, again information is scattered in time. Finally, in the frequency domain, FreqRISE puts most relevance on the fundamental frequency, while LRP places most relevance on the second harmonic of the fundamental frequency.
(a) Relevance maps in the STDFT domain.
(b) Relevance maps in time and frequency domains.
Figure 7: AudioMNIST: Digit task. Relevance maps in all three domains for a spoken digit 9 computed using LRP and FreqRISE.
(a) Relevance maps in the STDFT domain.
(b) Relevance maps in time and frequency domains.
Figure 8: AudioMNIST: Digit task. Relevance maps in all three domains for a spoken digit 0 computed using LRP and FreqRISE.
(a) Relevance maps in the STDFT domain.
(b) Relevance maps in time and frequency domains.
Figure 9: AudioMNIST: Gender task. Relevance maps in all three domains for a female speaker computed using LRP and FreqRISE.
(a) Relevance maps in the STDFT domain.
(b) Relevance maps in time and frequency domains.
Figure 10: AudioMNIST: Gender task. Relevance maps in all three domains for a male speaker computed using LRP and FreqRISE.
A.4 Results on synthetic data